# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ C.6. Green Stance Detection
Morio and Manning (2023) Most-frequent 46.7% 52.6% 36.8% 51.8% 25.6% 19.8% 41.2% 19.6% 17.5%
Morio and Manning (2023) Linear 66.0% 61.9% 50.3% 71.4% 44.5% 36.1% 52.0% 31.2% 27.0%
Morio and Manning (2023) BERT-base 71.0% 63.5% 51.6% 73.6% 48.1% 37.2% 50.2% 31.9% 25.8%
Morio and Manning (2023) ClimateBERT 71.8% 64.0% 52.8% 74.4% 48.9% 39.0% 50.2% 32.2% 26.8%
Morio and Manning (2023) RoBERTa-base 71.6% 64.5% 53.1% 73.8% 49.6% 38.3% 50.4% 33.4% 26.6%
Morio and Manning (2023) Longformer-base 73.7% 66.9% 54.6% 75.9% 53.0% 40.8% 52.5% 36.1% 28.6%
Morio and Manning (2023) Longformer-large 73.9% 68.8% 57.3% 76.5% 55.0% 44.1% 53.6% 38.7% 31.5%
(a) Custom metrics on LobbyMap (Morio and Manning, 2023).
Table 20. Reported performance for the datasets on stance detection (Part 2)
C.7. Question Answering
Source embeddings Quest. Def. Concepts Gen. inf_3 inf_all
Schimanski et al. (2024b) random 0.037 0.037 0.037 0.037 0.037 0.037
Schimanski et al. (2024b) BM25 0.113 0.114 0.126 0.139 0.172 0.174
Schimanski et al. (2024b) ColBERTv2 0.109 0.094 0.112 0.137 0.130 0.137
Schimanski et al. (2024b) DRAGON+ 0.139 0.121 0.161 0.160 0.161 0.160
Schimanski et al. (2024b) GTE-base 0.161 0.153 0.171 0.154 0.171 0.174
Schimanski et al. (2024b) text-embedding-ada-002 0.163 0.143 0.152 0.163 0.176 0.176
Schimanski et al. (2024b) text-embedding-3-small 0.163 0.143 0.152 0.163 0.179 0.179
Schimanski et al. (2024b) text-embedding-3-large 0.167 0.143 0.152 0.163 0.179 0.179
(a) F1-score (custom) of the different embedding models and information retrieval approaches (question, definition, concepts, generic, inf_3, inf_all) (Schimanski et al., 2024b)
Test In-house annotation
Source Model F1 BLEU METEOR F1 BLEU METEOR
Rony et al. (2022) ALBERT 81.6% 0.678 0.808 66.1% 0.416 0.694
(b) F1-score, BLEU and METEOR on CCMRC (Rony et al., 2022)
Source Model F1
Luccioni et al. (2020) large 85.5%
Luccioni et al. (2020) base 82%
(c) F1-scores (macro average) of RoBERTa on question answering (Luccioni et al., 2020).
Source Model F1
Spokoyny et al. (2023) Majority 4.11%
Spokoyny et al. (2023) Random 12.14%
Spokoyny et al. (2023) SVM 86.00%
Spokoyny et al. (2023) BERT 84.57%
Spokoyny et al. (2023) RoBERTa 85.61%
Spokoyny et al. (2023) DistilRoBERTa 84.38%
Spokoyny et al. (2023) Longformer 84.35%
Spokoyny et al. (2023) SciBERT 84.43%
Spokoyny et al. (2023) ClimateBERT 84.80%
(d) F1-scores (macro average) on ClimaINS Spokoyny et al. (2023)
Source Model Question Answer
R@1 R@5 R@1 R@5
Thulke et al. (2024) bge-base-en-v1.5 54.8 71.5 81.8 92.1
Thulke et al. (2024) bge-large-en-v1.5 55.8 73.6 83.3 93.1
Thulke et al. (2024) gtr-t5-large 48.8 67.4 79.6 90.1
Thulke et al. (2024) gtr-t5-xxl 47.6 66.3 79.2 89.7
Thulke et al. (2024) gte-large 50.7 68.2 80.9 91.4
Thulke et al. (2024) ember-v1 49.5 68.6 79.7 91.1
Thulke et al. (2024) instructor-large 50.0 68.2 81.7 91.8
Thulke et al. (2024) instructor-xl 53.3 69.7 83.3 92.1
(e) Recall (R@1 and R@5) of retrieving the correct document given the question or the answer from the IFT dataset (Thulke et al., 2024).
Model CITIES STATES CORP
@@ C.7. Question Answering
Finetuned on CDP (In-Domain)
@@
(f) MRR@10 scores on the three subsets of CLIMA-QA (Spokoyny et al., 2023). Models finetuned and evaluated on the same subset fall under In-Domain.
Task Raw Optimal Confidence
Net-zero target year 0.95 0.95 0.97
Reduction target year 0.84 0.90 0.90
Reduction base year 0.68 0.80 0.82
Reduction percentage 0.88 0.91 0.93
(g) Accuracy of the Q&A approach in assessing the quantitative properties of the net zero and reduction targets (Schimanski et al., 2023a)
Table 21. Reported performances for the datasets on Question/Answering and Retrieval tasks.
@@
Acc. Acc. F1 (macro)
Thulke et al. (2024) Stability-3B 71.4% 48.7% 52.6% 62.8%
Thulke et al. (2024) Pythia-6.9B 63.6% 22.9% 29.9% 48.9%
Thulke et al. (2024) Falcon-7B 62.9% 19.8% 39.9% 48.3%
Thulke et al. (2024) Mistral-7B 73.1% 19.0% 63.7% 53.7%
Thulke et al. (2024) Llama-2-7B 68.5% 51.1% 59.4% 62.6%
Thulke et al. (2024) Jais-13B 66.9% 26.4% 54.2% 54.4%
Thulke et al. (2024) Jais-13B-Chat 65.8% 66.4% 61.3% 64.6%
Thulke et al. (2024) Llama-2-Chat-7B 67.8% 72.0% 64.3% 68.5%
Thulke et al. (2024) Llama-2-Chat-13B 68.6% 79.3% 68.6% 71.4%
Thulke et al. (2024) Llama-2-Chat-70B 72.7% 88.2% 72.5% 77.8%
Thulke et al. (2024) ClimateGPT-7B 75.3% 86.6% 65.9% 77.1%
Thulke et al. (2024) ClimateGPT-13B 75.0% 89.0% 70.0% 78.0%
Thulke et al. (2024) ClimateGPT-70B 72.4% 89.9% 72.5% 77.7%
Thulke et al. (2024) ClimateGPT-FSC-7B 59.3% 17.2% 45.1% 46.2%
Thulke et al. (2024) ClimateGPT-FSG-7B 53.1% 17.4% 41.5% 42.1%
(a) F1 scores (macro average) and Accuracy on multiple climate benchmarks (Thulke et al., 2024).
Source Model F1
Spokoyny et al. (2023) Majority 20.10%
Spokoyny et al. (2023) Random 24.09%
Spokoyny et al. (2023) SVM -%
Spokoyny et al. (2023) BERT 70.57%
Spokoyny et al. (2023) RoBERTa 71.14%
Spokoyny et al. (2023) DistilRoBERTa 69.27%
Spokoyny et al. (2023) Longformer 67.72%
Spokoyny et al. (2023) SciBERT 68.45%
Spokoyny et al. (2023) ClimateBERT 69.44%
(b) F1-scores (macro average) on ClimaBench (Spokoyny et al., 2023)
Source Model Acc.
Vaghefi et al. (2023) hybrid ChatClimate 4.38
Vaghefi et al. (2023) ChatClimate 4.15
Vaghefi et al. (2023) GPT-4 2.62
(c) Accuracy on a set of 10 questions, human annotations (Vaghefi et al., 2023).
Table 22. Reported performance for the datasets on Question/Answering (few-shot and RAG) Tasks.
C.8. Deceptive Technique
Source Model F1
Trained on Logic applied on LogicCLimate
Jin et al. (2022) Electra-SA 22.72%
Jin et al. (2022) Electra 27.23%
@@ C.8. Deceptive Technique
Jin et al. (2022) Electra-SA 23.71%
Jin et al. (2022) Electra 29.37%
Baseline: Trained on Logic applied on Logic
Jin et al. (2022) Electra-SA 58.77%
Jin et al. (2022) Electra 53.31%
Jin et al. (2022) BERT 45.80%
(a) F1-scores (micro average) on fallacy detection (Jin et al., 2022). Electra-StructureAware is abbreviated Electra-SA
Source Model F1
Bhatia et al. (2021) BERT 59%
Bhatia et al. (2021) BERT* 62%
Bhatia et al. (2021) MTEXT 66%
Bhatia et al. (2021) MTEXT* 67%
Bhatia et al. (2021) MTEXTmulti 68%
Bhatia et al. (2021) MTEXT*multi 67%
Bhatia et al. (2021) SVM 49%
Bhatia et al. (2021) Human 70%
(b) F1-score (micro average) on neutralization techniques classification (Bhatia et al., 2021)
Source Model F1 (macro)
Coan et al. (2021) Logistic (unweighted) 68%
Coan et al. (2021) Logistic (weighted) 72%
Coan et al. (2021) SVM (unweighted) 66%
Coan et al. (2021) SVM (weighted) 72%
Coan et al. (2021) ULMFiT 72%
Coan et al. (2021) ULMFiT (weighted) 65%
Coan et al. (2021) ULMFiT (over sample) 55%
Coan et al. (2021) ULMFiT (focal Loss) 61%
Coan et al. (2021) ULMFiT-logistic 75%
Coan et al. (2021) ULMFiT-SVM 71%
Coan et al. (2021) RoBERTa 77%
Coan et al. (2021) RoBERTa-logistic 79%
Coan et al. (2021) Coder (Human) blAcc (macro): 87%
Thulke et al. (2024) Llama-2-Chat-70Bbla F1 (Bin.): 72.5%
Thulke et al. (2024) ClimateGPT-70Bblah F1 (Bin.): 72.5%
(c) F1-score (macro average and Binary) and Accuracy (macro average) on contrarian claim classification (Coan et al., 2021).
Table 23. Deceptive techniques
C.9. Environmental Performance Prediction
@@
Lin et al. (2023b) Baseline 1.173
Lin et al. (2023b) Lasso 0.848
Lin et al. (2023b) Gradient Boosting 0.818
Lin et al. (2023b) NOCATE(RB) 0.776
Lin et al. (2023b) NOCATE(CB) 0.763
Lin et al. (2023b) NOCATE(DB) 0.756
Lin et al. (2023b) SUSTAINABLESIGNALS(RB) 0.762
Lin et al. (2023b) SUSTAINABLESIGNALS(CB) 0.753
Lin et al. (2023b) SUSTAINABLESIGNALS(DB) 0.736
(a) MSE on Finch Score prediction (Finch Score range between 0 and 10) (Lin et al., 2023b)
Source Model F1
Clarkson et al. (2020) Random Forest 89.48%
Clarkson et al. (2020) XGBoost 95.73%
(b) F1-score (macro average) on good/bad CSR performer classification (Clarkson et al., 2020).
@@
(c) RMSE and wMAPE on ESG score prediction (ESG score range between 0 and 100) (Bronzini et al., 2023). We added the average RMSE and wMAPE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=50, scale=20) with the average value as prediction.
@@
Sharma et al. (2024) CLIP+MLP 0.54 0.88
Sharma et al. (2024) Sent.-BERT+MLP 0.57 0.91
@@
(d) MAE and RMSE on ESG unmanaged risk score prediction (ESG risk between 0 and 13.5) (Sharma et al., 2024). We added the average RMSE and MAE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=7, scale=2) with the average value a prediction.
@@
Mehra et al. (2022) majority class 0.5791 0.5682
Mehra et al. (2022) BERT 0.5985 0.4317
Mehra et al. (2022) ESGBERT 0.6709 0.793
(e) Accuracy on ESG-risk change classification (Mehra et al., 2022)
Table 24. Environmental Performance Prediction

### v2

@@ Appendix C Reported Performances
morio2023an Most-frequent 46.7% 52.6% 36.8% 51.8% 25.6% 19.8% 41.2% 19.6% 17.5%
morio2023an Linear 66.0 % 61.9 % 50.3 % 71.4 % 44.5 % 36.1 % 52.0 % 31.2 % 27.0 %
morio2023an BERT-base 71.0 % 63.5 % 51.6 % 73.6 % 48.1 % 37.2 % 50.2 % 31.9 % 25.8 %
morio2023an ClimateBERT 71.8 % 64.0 % 52.8 % 74.4 % 48.9 % 39.0 % 50.2 % 32.2 % 26.8 %
morio2023an RoBERTa-base 71.6 % 64.5 % 53.1 % 73.8 % 49.6 % 38.3 % 50.4 % 33.4 % 26.6 %
morio2023an Longformer-base 73.7 % 66.9 % 54.6 % 75.9 % 53.0 % 40.8 % 52.5 % 36.1 % 28.6 %
morio2023an Longformer-large 73.9 % 68.8 % 57.3 % 76.5 % 55.0 % 44.1 % 53.6 % 38.7 % 31.5 %
(e) Performance on LobbyMap [morio2023an] using custom metrics described in the study.
Table 20: Reported performance for the datasets on stance detection
Question Answering
embeddings question definition concepts generic inf_3 inf_all
random 0.037 0.037 0.037 0.037 0.037 0.037
BM25 0.113 0.114 0.126 0.139 0.172 0.174
ColBERTv2 0.109 0.094 0.112 0.137 0.130 0.137
DRAGON+ 0.139 0.121 0.161 0.160 0.161 0.160
GTE-base 0.161 0.153 0.171 0.154 0.171 0.174
text-embedding-ada-002 0.163 0.143 0.152 0.163 0.176 0.176
text-embedding-3-small 0.163 0.143 0.152 0.163 0.179 0.179
text-embedding-3-large 0.167 0.143 0.152 0.163 0.179 0.179
(a) Results for the F1-score (custom) of the different embedding models and information retrieval approaches (question, definition, concepts, generic, inf_3, inf_all) aggregated across all top-k values (5, 10, 15). [schimanski-etal-2024-climretrieve]
Source Model Test In-house annotation test
F1-score BLEU METEOR F1-score BLEU METEOR
cliamtebot_2022 ALBERT 81.6 % 0.678 0.808 66.1 % 0.416 0.694
(b) Performance on CCMRC cliamtebot_2022
Task Data Accuracy (% of data)
net zero target year raw/optimal 0.95 (1.0)
confidence-tuned 0.97 (0.5)
reduction target year raw 0.84 (1.0)
optimal 0.90 (0.97)
confidence-tuned 0.90 (0.8)
reduction base year raw 0.68 (1.0)
optimal 0.80 (0.84)
confidence-tuned 0.82 (0.62)
reduction percentage raw 0.88 (1.0)
optimal 0.91 (0.96)
confidence-tuned 0.93 (0.35)
(c) Accuracy of the Q&A approach in assessing the quantitative properties of the net zero and reduction targets [tobias_schimanski_climatebert-netzero_2023]
Source Model F1 (macro)
spokoyny2023answering Majority 4.11 %
spokoyny2023answering Random 12.14 %
spokoyny2023answering SVM 86.00 %
spokoyny2023answering BERT 84.57 %
spokoyny2023answering RoBERTa 85.61 %
spokoyny2023answering DistilRoBERTa 84.38 %
spokoyny2023answering Longformer 84.35 %
spokoyny2023answering SciBERT 84.43 %
spokoyny2023answering ClimateBERT 84.80 %
(d) Performance on ClimaINS spokoyny2023answering
CDP-CITIES CDP-STATES CDP-CORP
Model MRR@10 MRR@10 MRR@10
@@
Finetuned on CDP
In-Domain In-Domain In-Domain
@@
(e) MRR@10 scores for BM25, ClimateBERT and MSMARCO-MiniLM on the three subsets of CLIMA-QA. Models finetuned and evaluated on same subset fall under In-Domain.
Source Model F1-score (macro)
luccioni_analyzing_2020 RoBERTa-large 85.5 %
luccioni_analyzing_2020 RoBERTa-base 82 %
(f) Performance on question answering [luccioni_analyzing_2020] (2 labels: answer, not) mapping a sentence to one of the TCFD questions
Table 21: Reported performance for the datasets on Question/Answering (Retrieval) Tasks
Source Model Accuracy
s_vaghefi_chatclimate_2023 hybrid ChatClimate 4.38
s_vaghefi_chatclimate_2023 ChatClimate 4.15
s_vaghefi_chatclimate_2023 GPT-4 2.62
(a) ”Accuracy” on a set of 10 questions, human annotations s_vaghefi_chatclimate_2023
Source Model F1 (macro)
spokoyny2023answering Majority 20.10%
spokoyny2023answering Random 24.09%
spokoyny2023answering SVM - %
spokoyny2023answering BERT 70.57%
spokoyny2023answering RoBERTa 71.14%
spokoyny2023answering DistilRoBERTa 69.27%
spokoyny2023answering Longformer 67.72%
spokoyny2023answering SciBERT 68.45%
spokoyny2023answering ClimateBERT 69.44%
(b) Performance on ClimaBench spokoyny2023answering
@@
F1 (macro) F1 (unspecified) F1 (macro)
thulke2024climategpt Stability-3B 71.4 48.7 52.6 62.8
thulke2024climategpt Pythia-6.9B 63.6 22.9 29.9 48.9
thulke2024climategpt Falcon-7B 62.9 19.8 39.9 48.3
thulke2024climategpt Mistral-7B 73.1 19.0 63.7 53.7
thulke2024climategpt Llama-2-7B 68.5 51.1 59.4 62.6
thulke2024climategpt Jais-13B 66.9 26.4 54.2 54.4
thulke2024climategpt Jais-13B-Chat 65.8 66.4 61.3 64.6
thulke2024climategpt Llama-2-Chat-7B 67.8 72.0 64.3 68.5
thulke2024climategpt Llama-2-Chat-13B 68.6 79.3 68.6 71.4
thulke2024climategpt Llama-2-Chat-70B 72.7 88.2 72.5 77.8
thulke2024climategpt ClimateGPT-7B 75.3 86.6 65.9 77.1
thulke2024climategpt ClimateGPT-13B 75.0 89.0 70.0 78.0
thulke2024climategpt ClimateGPT-70B 72.4 89.9 72.5 77.7
thulke2024climategpt ClimateGPT-FSC-7B 59.3 17.2 45.1 46.2
thulke2024climategpt ClimateGPT-FSG-7B 53.1 17.4 41.5 42.1
(c) Results on climate benchmarks[thulke2024climategpt]
Table 22: Reported performance for the datasets on Question/Answering (few-shot and RAG) Tasks.
Deceptive Technique
Source Model F1 (micro)
Trained on general domain
jin-etal-2022-logical Electra-StructureAware 22.72 %
jin-etal-2022-logical Electra 27.23 %
@@
jin-etal-2022-logical Electra-StructureAware 23.71 %
jin-etal-2022-logical Electra 29.37 %
Baseline: Performance on Logic
jin-etal-2022-logical Electra-StructureAware 58.77 %
jin-etal-2022-logical Electra 53.31 %
jin-etal-2022-logical BERT 45.80 %
(a) Performances on jin-etal-2022-logical fallacy detection (13 labels)
Source Model F1 (micro)
bhatia_automatic_2021-1 BERT 59 %
bhatia_automatic_2021-1 BERT* 62 %
bhatia_automatic_2021-1 MTEXT 66 %
bhatia_automatic_2021-1 MTEXT* 67 %
bhatia_automatic_2021-1 MTEXTmulti 68 %
bhatia_automatic_2021-1 MTEXT*multi 67 %
bhatia_automatic_2021-1 SVM 49 %
bhatia_automatic_2021-1 Human 70 %
(b) Performance on bhatia_automatic_2021-1
Source Model F1-score (macro)
coanComputerassistedClassificationContrarian2021 Logistic (unweighted) 68 %
coanComputerassistedClassificationContrarian2021 Logistic (weighted) 72 %
coanComputerassistedClassificationContrarian2021 SVM (unweighted) 66 %
coanComputerassistedClassificationContrarian2021 SVM (weighted) 72 %
coanComputerassistedClassificationContrarian2021 ULMFiT 72 %
coanComputerassistedClassificationContrarian2021 ULMFiT (weighted) 65 %
coanComputerassistedClassificationContrarian2021 ULMFiT (over sample) 55 %
coanComputerassistedClassificationContrarian2021 ULMFiT (focal Loss) 61 %
coanComputerassistedClassificationContrarian2021 ULMFiT-logistic 75 %
coanComputerassistedClassificationContrarian2021 ULMFiT-SVM 71 %
coanComputerassistedClassificationContrarian2021 RoBERTa 77 %
coanComputerassistedClassificationContrarian2021 RoBERTa-logistic 79 %
Accuracy (macro)
coanComputerassistedClassificationContrarian2021 Coder (Human) 87 %
Binary Classification
[thulke2024climategpt] Llama-2-Chat-70B 72.5 %
[thulke2024climategpt] ClimateGPT-70B 72.5 %
(c) Performance on contrarian claim classification [coanComputerassistedClassificationContrarian2021]
Table 23: Deceptive techniques
Environmental Performance Prediction
@@
linSUSTAINABLESIGNALSijcai2023 Baseline 1.173
linSUSTAINABLESIGNALSijcai2023 Lasso 0.848
linSUSTAINABLESIGNALSijcai2023 Gradient Boosting 0.818
linSUSTAINABLESIGNALSijcai2023 NOCATE(RB) 0.776
linSUSTAINABLESIGNALSijcai2023 NOCATE(CB) 0.763
linSUSTAINABLESIGNALSijcai2023 NOCATE(DB) 0.756
linSUSTAINABLESIGNALSijcai2023 SUSTAINABLESIGNALS(RB) 0.762
linSUSTAINABLESIGNALSijcai2023 SUSTAINABLESIGNALS(CB) 0.753
linSUSTAINABLESIGNALSijcai2023 SUSTAINABLESIGNALS(DB) 0.736
(a) Performance on the task of Finch Score prediction (Finch Score range between 0 and 10) linSUSTAINABLESIGNALSijcai2023
@@
(b) Performance on ESG score prediction (ESG score range between 0 and 100) bronzini_glitter_2023. We added the average RMSE and wMAPE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=50, scale=20) with the average value a prediction.
Source Model F1 (macro)
clarkson_nlp_us_csr Random Forest 89.48 %
clarkson_nlp_us_csr XGBoost 95.73 %
(c) clarkson_nlp_us_csr (2 classes: good performer, bad performer)
@@
Greenscreen CLIP+MLP 0.54 0.88
Greenscreen Sentence-BERT+MLP 0.57 0.91
@@
(d) Performance on ESG unmanaged risk score prediction from Greenscreen (ESG risk between 0 and 13.5). We added the average RMSE and MAE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=7, scale=2) with the average value a prediction.
@@
Mehra_2022 majority class 0.5791 0.5682
Mehra_2022 BERT 0.5985 0.4317
Mehra_2022 ESGBERT 0.6709 0.793
(e) Performance on Mehra_2022’s tasks of ESG risk change classification
Table 24: Environmental Performance Prediction